OpenStudy (anonymous):

The ages of Edna,Ellie,& Elsa areconsecutive integers. The sum of their ages is 108. What are the ages,youngest to oldest?

OpenStudy (agent47):

x+(x+1)+(x+2)=108 3x+3=108 3(x+1)=108

OpenStudy (agent47):

x+1=36 x=35

OpenStudy (agent47):

Their ages are: 35, 36, 37

OpenStudy (anonymous):

Thanks so much. You made me feel dumb--BUT I got it right :)